HIGH RESOLUTION LASER EXCITATION SPECTROSCOPY OF BARIUM MONOSULFIDE

\underline{G. LI}, J.-G. WANG, and P. F. BERNATH, Department of Chemistry, University of York, Heslington, York, YO10 5DD, UK.

BaS molecules were synthesized using a Broida-type oven and detected by laser excitation spectroscopy using a single mode Ti:Sapphire laser. High resolution spectra of BaS were recorded in the 12200−12765 cm^{−1} region. The BaS spectra contain several bands, and preliminary assignments and least-squares fits were carried out using ground state microwave data. The main bands are provisionally assigned to the A^{1}Π − X^{1}Σ^{+} transition of BaS. Further analysis of the minor isotopologues will be performed in order to obtain a secure vibrational assignment of the upper state levels. More laser scans will also be carried out from 12200 cm^{−1} to the lower frequency limit of the Ti:Sapphire laser. Preliminary results on this ongoing project will be presented.